# A NIRCam-dark galaxy detected with the MIRI/F1000W filter in the MIDIS/JADES Hubble Ultra Deep Field

### v1

@@ 5 Summary and conclusions
A low-redshift solution, , would imply that Cerberus is a dusty galaxy with strong emission from polycyclic aromatic hydrocarbons or warm dust heated by an active galactic nucleus. The low mass obtained for this solution, M, would point to an unknown type of dwarf galaxy with large dust content and no PAH depletion.
The second possibility is a dusty starburst or post-starburst galaxy, or a galaxy hosting a mid-infrared bright obscured active galactic nucleus, at with a stellar mass M and around mag of extinction.
The third possibility is a galaxy, the colors of which could be understood if we assume its spectral energy distribution is dominated by an obscured active nucleus with a bright torus significantly contributing to the flux in the rest-frame optical range, and the super-massive black hole to stellar mass ratio is around 0.2. This fit would leave room for some additional contribution from stellar emission, perhaps dominating the rest-frame ultraviolet, and providing a stellar mass around M. The latter number is, however, very uncertain due to the overshining effect of the torus. Alternatively, the spectral energy distribution can be reproduced with a relatively massive ( M) dusty ( mag) galaxy.

### v2

@@ 5 Summary and conclusions
A low-redshift solution, , would imply that Cerberus is a dusty galaxy with strong emission from polycyclic aromatic hydrocarbons or warm dust heated by an active galactic nucleus. The low mass obtained for this solution, M⊙, would point to an unknown type of dwarf galaxy with large dust content and no PAH depletion.
The second possibility is a dusty starburst or post-starburst galaxy, or a galaxy hosting a mid-infrared bright obscured active galactic nucleus, at with a stellar mass M⊙ and high extinction.
The third possibility is a galaxy, the red color emanating from a spectral energy distribution dominated by emission from an obscured active nucleus with a bright torus significantly contributing to the flux in the rest-frame optical range, with shorter wavelengths dominated by a M⊙ galaxy.
@@
Appendix A Fitting codes used to analyze Cerberus spectral energy distribution
In this Appendix, we give details of the codes used to analyze the SED of Cerberus and obtain its possible redshifts if it were an extragalactic object, as well as the implications of those redshifts in its physical properties.
We first used the eazy code (Brammer et al., 2008) to fit the data in 2 ways: one using actual (low ) flux measurements for all filters, and another replacing all measurements by upper limits that templates were not allowed to surpass (achieved with a modified version of the code). We employed in the fits all v1.3 templates for stellar-dominated galaxies 111As listed here: https://github.com/gbrammer/eazy-photoz/tree/master/templates., which include a dusty galaxy with a high-equivalent width emission-line spectrum. We also allowed combinations of stellar-only templates with the new models for little red dots and high-redshift AGNtorus recently added based on JWST data (Killi et al., 2023). We did not impose any prior and worked with minimum photometric redshift estimates in the range .
The synthesizer-AGN code assumes that the SED can be modeled with a composite stellar population (Pérez-González et al., 2003, 2008) and AGN emission coming from the accretion disk and the dust torus (Pérez-González et al., 2024). The stellar emission includes a young and a more evolved star formation event, each one described by a delayed exponential function with timescales between 1 Myr and 1 Gyr, and with ages from 1 Myr up to the age of the Universe at the redshift of the source. The attenuations of the emission from each stellar population are independent and described by a Calzetti et al. (2000) law, with AV values ranging from 0 to 10 mag for each population. The stellar emission is described by the Bruzual & Charlot (2003) models, assuming a Chabrier (2003) IMF with stellar mass limits between 0.1 and 100 , and the nebular emission is also considered (Pérez-González et al. 2003). The AGN emission is modeled with a QSO average spectrum (Vanden Berk et al., 2001; Glikman et al., 2006). The dust emission from the AGN is modeled with the self-consistent templates of AGN tori presented in Siebenmorgen et al. (2015), and dust emission linked to star formation is added using the models from Draine & Li (2007).
CIGALE (Boquien et al., 2019) SED-fitting analysis assume an SFH modelled using a constant star-formation rate with ages ranging from 1 to 100 Myr. We adopted the stellar populations models from Bruzual & Charlot (2003) with solar metallicity, and the Chabrier (2003) IMF. We included nebular continuum and emission lines using solar metallicity, electron density of 100 cm-3, and ionised parameter equal to =2. The dust attenuation and far-IR emission uses the Calzetti Law (Calzetti et al., 2000) and the Draine et al. (2014) models, respectively. AGN emission is added using the Fritz et al. (2006) models following the initial parameters suggested by Ciesla et al. (2015).
The obtained photometric redshifts solutions are: (provided by prospector and LePHARE++), (very prominent solution for by eazy, also detected weakly by prospector and LePHARE++), (obtained by all codes), (obtained by LePHARE++), and (obtained by bagpipes and LePHARE++). We note that the solution corresponds to a PAH line entering the F1000W passband, the solution would imply Pa covered by the filter, for , H would contribute to the MIRI-F1000W flux, and means that the F1000W emission could be enhanced by H and/or [O III]4960,5008 emission.
We fitted the SED shown in Figure 4, which includes all flux data points with and upper limits for the rest of bands, using the codes described above. We checked that our results concerning the photometric redshift and physical properties (for the two parameters mentioned in the main text, the stellar mass and the dust attenuation) do not change significantly if we only use fluxes and upper limits, or just fluxes. More specifically, the photometric redshift solutions are the same, but with different statistical weights (i.e., the peaks vary their relative strength). Concerning the physical properties, variations were very similar to the systematic differences between the distinct estimations for the two properties that are consistent among all codes and techniques, given that they are constrained by the two robust observational properties of Cerberus, i.e., the low luminosity and red color. More specifically, for the stellar mass we obtained variations of 0.3 dex when using higher signal-to-noise data for and , and 0.7 dex for . For the dust content, we obtained differences of mag in all cases, always remaining relatively high, above 3 mag, or finding a similarly obscured torus model.